Two point charges are fixed on the y axis: a negative point charge q1 = -26 µC at y1 = +0.20 m and a positive point charge q2 at y2 = +0.30 m. A third point charge q = +8.6 µC is fixed at the origin. The net electrostatic force exerted on the charge q by the other two charges has a magnitude of 22 N and points in the +y direction. Determine the magnitude of q2.
Answer #1

q2 q1-26uC 0.2m F1 F2 Net F-F1-F2-22N 0.3m q-8.6uC

F_{1}-F_{2}=22N

kqq_{1}/r_{1}^{2}-kqq_{2}/r_{2}^{2}=22N

9*10^{9}*8.6*10^{-6}C*2.6*10^{-5}C/(0.2m)^{2}-kqq_{2}/r_{2}^{2}=22N

50.31-kqq_{2}/r_{2}^{2}=22N

kqq_{2}/r_{2}^{2}=28.31N

9*10^{9}*8.6*10^{-6}C*q_{2}*10^{-6}/(0.3m)^{2}=28.31N

0.86*q_{2}=28.31

ANSWER: q_{2}=32.91mu C

===================